The auto parts department of an automotive dealership sends out a mean of 3.3 special orders daily. What is the probability that, for any day, the number of special orders sent out will be more than 4? Round your answer to four decimal places.
Answer #1

Solution :

Given that ,

mean = = 3.3

Using poison probability formula,

P(X > 4) = 1 - [P(X = 0)- P(X = 1)- P(X = 2)- P(X = 3)- P(X = 4)]

= (e-3.3 * 3.30) / 0! + (e-3.3 * 3.31) / 1! + (e-3.3 * 3.32) / 2! + (e-3.3 * 3.33) / 3! + (e-3.3 * 3.34) / 4!

Probability = 0.2374